For the titration of 50. mL of 0.10 M ammonia with 0.10 M HCl, calculate the pH (a) before HCl is added, (b) after 20 ml of acid is added, and (c) after half of the NH3 has been neutralized
Answer #1

I´ll answer the first question, if you want the others answer, attach them in a nwe question.

a)

ammonia=0.1mol/L... in 50 mL=

C1/V1=C2/V2

C2=C1*V2/V1=0.1*0.05/1=5*10-3

pOH=-log5*10-3

pH=14-(-log5*10-3)=14-2.3=11.7